# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
Driven-dissipative many-body system supports nontrivial quantum phases absent in equilibrium. As a prominent example, the interplay between coherent driving and collective dissipation can lead to a dynamical quantum phase that spontaneously breaks time-translation symmetry. This so-called boundary time crystal (BTC) is fragile in the presence of local dissipation, which inevitably relaxes the system to a stationary state. In this work, we demonstrate a robust BTC that is intrinsically induced by local dissipation. We provide extensive numerical evidences to support existence of the BTC and study its behaviors in different regimes. In particular, with decreasing interaction range, we identify a transition from classical limit cycles to quantum BTCs featuring sizable spatial correlations. Our studies significantly broaden the scope of nonequilibrium phases and shed new light on experimental search for dynamical quantum matter.
@@
Introduction.—Spontaneous symmetry breaking in the temporal dimension manifests itself in a nontrivial quantum phase: the continuous time crystal (CTC) [1]. In the CTC phase, a small perturbation brings the system away from the stationary state and triggers a persistent, ordered oscillation of local observables [2, 3, 4, 5]. While existence of this intriguing phase in equilibrium has been questioned by a series of studies [6, 7, 8], it has been predicted that CTC can be realized in a driven-dissipative setup [9, 10, 11, 12, 13, 14, 15, 16, 17]. This possibility is highlighted by a number of recent experiments observing self-sustained oscillations, such as in ultracold atoms [18, 19, 20], thermal Rydberg gases [21, 22, 23, 24, 25, 26], and solid-state materials [27, 28].
@@
The dissipative version of the CTC is also called a boundary time crystal (BTC), considering the system as a boundary in contact with a bulk environment [29]. The prototype model of the BTC considers a system of spin-1/2 particles subjected to a superradiant decay [30], which can be induced by a propagating bath mode in the Markovian regime [see Fig. 1(a)]. However, a realistic setup inevitably carries local decay channels, which immediately destroy the established oscillation pattern in the long-time limit [31, 32], as shown in Fig. 1(b).
Inspired by recent experiments in Rydberg atomic ensembles [22, 23, 24] we study in this Letter a one-dimensional (1D) lattice model, where long-range interacting spin-1 particles experience strictly local decay [see Fig. 1(c)]. While local dissipations usually erase quantum correlations [33], we show that, remarkably, they can induce a BTC phase here [see Fig. 1(d)]. At the mean-field level, this BTC corresponds to a classical limit cycle, which originates from competition between distinct spin coherences. We perform extensive numerical simulations to investigate the quantum nature and origin of the BTC. In the case of all-to-all interactions, we show that the Liouvillian spectrum exhibit gapless and ordered features, while persistent oscillation of the two-time correlation function unambiguously reveals a spontaneous time-translation symmetry breaking. For a power-law interaction $\sim r^{-\alpha}$, our analysis reveals a critical $\alpha_{c}$, above which the interaction is too short-ranged to support the BTC. Interestingly, $\alpha_{c}$ is larger than the dimension of the system, giving rise to a quantum BTC without a classical correspondence in the regime $1<\alpha<\alpha_{c}$. In a quantum BTC, considerable spatial correlations can be rapidly established from an uncorrelated initial state.
@@
Figure 1: (a) and (b) show the schematic of the superradiance-based BTC and its mean-field evolution of the order parameter $m_{z}$, respectively. Here, an arbitrarily small local decay ($\gamma\neq 0$) can destroy the established oscillation. (c) and (d) show the proposed spin-1 BTC model and its mean-field evolution of the order parameter $s_{z}$, respectively. Here, the oscillation is robust against local dissipation and persists within a finite range of the decay rate $\gamma$.
@@
where the spin-1 operators are defined as $s_{i}^{x}=(|0\rangle_{i}\langle+|+|0\rangle_{i}\langle-|+\mathrm{H.c.})/\sqrt{2}$, $s_{i}^{z}=|+\rangle_{i}\langle+|-|-\rangle_{i}\langle-|$, and $n_{i}=(s_{i}^{z})^{2}=|+\rangle_{i}\langle+|+|-\rangle_{i}\langle-|$. In such a $\mathsf{V}$-type excitation scheme, the spin state |$\ket{0}$ is coherently coupled to states $\ket{+}$ and $\ket{-}$ (energetically separated by $E$) with a Rabi frequency $\Omega$ and a detuning $\Delta$ from the average energy level of the exited states. Meanwhile, spins in states $\ket{\pm}$ feature a power-law interaction $V_{ij}=C/|i-j|^{\alpha}$. In addition to coherent drivings and interactions, each spin in states $\ket{\pm}$ decays individually to the state |$\ket{0}$ in a Markovian manner with a rate $\gamma$. The system density matrix $\rho$ then evolves according to a Lindblad master equation $\partial_{t}\rho=\mathcal{L}(\rho)$, where the Liouvillian operator acts as
@@
Mean-field analysis.—We first treat the system at the mean-field level, in which we neglect the correlation between different sites and assume that local operators $o_{i}$ have a homogeneous expectation value, i.e., $\langle o_{i}o_{j}\rangle=\langle o_{i}\rangle\langle o_{j}\rangle$ for $j\neq k$ and $\langle{o}_{i}\rangle=\langle{o}_{j}\rangle$. This treatment will generate a set of nonlinear equations [37], whose nonlinear dynamics is solely determined by the collective interaction strength $\chi=\sum_{j}V_{ij}$.
We then perform stability analysis of the equations of motion [38] and study the long-time behavior of the system. First, when states $\ket{\pm}$ are degenerate ($E=0$), the system only has stable fixed points, corresponding to a stationary phase (SP) where observables rapidly decay to the steady-state values. Within the SP, by tuning the driving parameters $\Omega$ and $\Delta$, the system can undergo a first-order phase transition between a bistable phase and a trivial one. In fact, although states $\ket{\pm}$ decay independently, the mean-field equations at $E=0$ is equivalent to those of a spin-1/2 model, which is widely used to describe the optical bistability in Rydberg atomic ensembles [39, 40, 41, 42, 43].
As the energy separation $E$ increases, the competition between states $\ket{\pm}$ comes into play, which is quantified by the order parameter $s_{z}=\langle s_{i}^{z}\rangle$ measuring their population difference. Near the intermediate driving ($|\Delta|<\Omega$), the system supports a Hopf bifurcation at which the stability of a fixed point suddenly changes (see the Supplemental Material for details [37]). Figure 1(d) shows a typical case $E=4$, $\Delta=-1$, $\Omega=3$, $\chi=16$, $\gamma=1$ (used throughout this work), in which $s_{z}$ exhibits a persistent periodic oscillation, whose trajectory is depicted by the middle panel of Fig. 2(b). The oscillation pattern of observables is akin to the coupled pendulum model [44], with the two distinct atomic coherences acting as the harmonic oscillators. This is distinct from the limit cycle in purely dissipative cases, e.g., the Lotka–Volterra predator–prey model [45, 46, 47]. Such an understanding is confirmed by adiabatic elimination of coherence terms: in their absence, the system ceases to oscillate.
@@
We first study the Liouvillian spectrum $\{\lambda_{k}\}$ with $\mathcal{L}(\rho_{k})=\lambda_{k}\rho_{k}$ in the BTC phase. As the system size increases from 10 to 20, we identify multiple ordered branches of eigenvalues $\lambda_{k}$ shifting towards the imaginary axis, which is a universal feature for systems supporting limit cycles [52]. The inset of Fig. 3(a) depicts the low-lying spectrum for $N=20$, where we can identify a branch of eigenvalues evenly spaced along the imaginary-axis direction, as indicated by the blue dots close to the dashed grids $0,\pm\omega,\pm 2\omega,\cdots$, with $\omega$ the oscillation frequency of the limit cycle. This branch will collapse onto the imaginary axis in the thermodynamic limit, responsible for all possible oscillation patterns of the BTC. The real part of such a gapless branch follows an algebraic decay $\mathrm{Re}[\lambda_{k}]\propto-N^{-\beta_{k}}$. Importantly, smaller frequency components approach the imaginary axis with a higher rate $\beta_{k}$ [see Fig. 3(a) for the gap closing process of ±$\pm\omega$ and ±$\pm 2\omega$ components]. As a result, the gapless branch will gradually form a parabolic shape, responsible for the quantum state diffusion in the phase space [52].
@@
In addition to ordered branches, we find eigenvalues whose imaginary parts are incommensurate with $\omega$. The ones closest to the imaginary axis are indicated by the red diamonds in Fig. 3(a). These Liouvillian eigenstates are gapped, and determine the time scale at which a stable oscillation can be established. The gap $\approx 0.594\gamma$ drawn from an algebraic fitting shows a nice agreement with the decay rate $\approx 0.602\gamma$ toward the limit cycle.
@@
Figure 4: Scaling of the steady-state fluctuations $F^{2}$ with $N$. (a) and (b) show results for the BTC phase ($\Delta=-1$) and the SP ($\Delta=-7$), respectively. The insets show the time evolution of the fluctuations ($N=80$ for the BTC and $N=40$ for the SP), obtained from the MPS (red curves) and the CME (blue curves). The errorbars indicate the temporal fluctuations caused by finite trajectory realizations.
@@
In fact, a time crystal defined by Eq. (3) possesses not only temporal order but also long-range spatial order, which is indeed true for all-to-all interaction since $\mathcal{C}_{\mathrm{off}}\approx F^{2}/N=G(0)$ remains finite in the large-$N$ limit. As shown in Fig. 5(c), $\mathcal{C}_{\mathrm{off}}$ decreases monotonically with an increasing $\alpha$ and eventually vanishes when entering the SP  [^2] This trend is verified by the MPS for $N=20$ [37]. The critical property is then analyzed through finite-size scaling by assuming a universal form of $\mathcal{C}_{\mathrm{off}}=N^{\zeta/\nu}\mathcal{G}[N^{1/\nu}(\alpha-\alpha_{c})]$. We obtain a nice data collapse [see the inset of Fig. 5(c)] and extract a critical value $\alpha_{c}\approx 1.22$ together with two critical exponents $(\zeta,\nu)\approx(-0.56,1.40)$. We note that $\alpha_{c}$ is generally larger than 1 but not rigid under variation of system parameters, typical for long-range interacting systems [63].
@@
The proposed model can be potentially realized in artificial quantum systems, such as trapped ions or neutral atoms. In the former case, tunable power-law interactions with $30\leq\alpha<$ 3 can be achieved [66, 67, 68], aligning with the range considered in this work. For the latter, all-to-all interactions can be induced by cavity QED [69, 70] or Rydberg dressing in atomic ensembles [71]. Our model can also be generalized to higher dimensions, where the long-time dynamics may be accessed by improved Krylov methods [72]. In higher dimensions, dipolar ($\alpha=3$) or van der Waals ($\alpha=6$) interactions offered by Rydberg-atom arrays [73, 74, 75] may be sufficiently long-ranged to stabilize the BTC. In addition, our studies may reveal a new universality class for nonequilibrium phases [76], and open the door to exploring versatile dynamical phases, such as multistability [77, 78, 79] and measurement-induced phase transition [80].
@@ Acknowledgements.
Acknowledgments.—We thank T. Pohl, X.L. Li, Y. Zhang, G.M. Zhang for providing helpful suggestions, and X.Y. Wen, Z. Fu, C.P. Qin for sharing computation resources. F.Y. acknowledge valuable discussions with T. Vovk, L. Sieberer, and H. Pichler. This work is supported by the Innovation Program for Quantum Science and Technology (2021ZD0302100) and by the NSFC (Grants No. 12361131576, No. 92265205). B.B. acknowledges the support from funding by the French National Research Agency (ANR) under project ANR-24-CPJ1-0150-01 and a research grant (42085) from VILLUM FONDEN. K.M. acknowledges the support from the Carlsberg Semper Ardens project QCooL. Work in Innsbruck is supported by the ERC Starting grant QARA (Grant No. 101041435). Note added: During preparation of this manuscript, we became aware of a related work on three-level dissipative time crystals induced by a dipolar interaction in 2D [81].

### v2

@@ Abstract
Driven-dissipative many-body system supports nontrivial quantum phases absent in equilibrium. As a prominent example, the interplay between coherent driving and collective dissipation can lead to a dynamical quantum phase that spontaneously breaks time-translation symmetry. This so-called boundary time crystal (BTC) is fragile in the presence of local dissipation, which can easily relax the system to a stationary state. In this work, we demonstrate a robust BTC that is intrinsically induced by local dissipation. We provide extensive numerical evidences to support existence of the BTC and study its behaviors in different regimes. In particular, with decreasing interaction range, we identify a transition from classical limit cycles to quantum BTCs featuring sizable spatial correlations. Our studies significantly broaden the scope of nonequilibrium phases and shed new light on experimental search for dynamical quantum matter.
@@
Introduction.—Spontaneous symmetry breaking in the temporal dimension manifests itself in a nontrivial quantum phase: the continuous time crystal (CTC) [1]. In the CTC phase, a small perturbation brings the system away from the stationary state and triggers a persistent, ordered oscillation of local observables [2, 3, 4, 5, 6]. While existence of this intriguing phase in equilibrium has been questioned by a series of studies [7, 8, 9], it has been predicted that CTC can be realized in a driven-dissipative setup [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. This possibility is highlighted by a number of recent experiments observing self-sustained oscillations, such as in ultracold atoms [20, 21, 22], thermal Rydberg gases [23, 24, 25, 26, 27, 28], and solid-state materials [29, 30].
@@
The dissipative version of the CTC is also called a boundary time crystal (BTC), considering the system as a boundary in contact with a bulk environment [31]. The prototype model of the BTC considers a system of spin-1/2 particles subjected to a superradiant decay [32], which can be induced by a common Markovian bath [see Fig. 1(a)]. However, a realistic setup inevitably carries local decay channels, which can easily destroy the established oscillation [33, 34, 35], as exemplified in Fig. 1(b). The robustness of the BTC can be enhanced by a multi-level setting [36], but a local decay generally breaks the underlying $\mathrm{SU}(n)$ algebra and destabilizes the BTC.
@@
Figure 1: (a) and (b) show the schematic of the superradiance-based BTC and its mean-field evolution of the magnetization $m_{z}$, respectively. Here, a small local decay $\gamma=0.01\kappa$ (with $\kappa$ the collective decay rate and $\Omega=1.5\kappa$) can destroy the oscillation. (c) and (d) show the proposed spin-1 BTC model and its mean-field evolution of the order parameter $s_{z}$, respectively. Here, the oscillation is induced solely by local dissipation and persists within a finite range of the decay rate $\gamma$.
In this Letter, we investigate whether a BTC can be induced solely by local dissipation. The considered model is a one-dimensional (1D) spin chain, where long-range interacting spin-1 particles experience strictly local decay [see Fig. 1(c)]. While local dissipation usually erases quantum correlations [37], we show that, remarkably, they can induce a BTC here [see Fig. 1(d)]. At the mean-field level, this BTC corresponds to a classical limit cycle, which originates from competition between distinct spin excitations. We perform extensive numerical simulations to investigate the quantum nature and origin of the BTC. In the case of all-to-all interactions, we show that the Liouvillian spectrum exhibit gapless and ordered features, while persistent oscillation of the two-time correlation function unambiguously reveals a spontaneous time-translation symmetry breaking. For a power-law interaction $\sim r^{-\alpha}$, our analysis reveals a critical $\alpha_{c}$, above which the interaction is too short-ranged to support the BTC. Interestingly, $\alpha_{c}$ is larger than the dimension of the system, giving rise to a quantum BTC without a classical correspondence in the regime $1<\alpha<\alpha_{c}$. In a quantum BTC, considerable spatial correlations can be rapidly established from an uncorrelated initial state.
@@
where the spin-1 operators are defined as $s_{i}^{x}=(|0\rangle_{i}\langle+|+|0\rangle_{i}\langle-|+\mathrm{H.c.})/\sqrt{2}$, $s_{i}^{z}=|+\rangle_{i}\langle+|-|-\rangle_{i}\langle-|$, and $n_{i}=(s_{i}^{z})^{2}=|+\rangle_{i}\langle+|+|-\rangle_{i}\langle-|$. In such a $\mathsf{V}$-type excitation scheme, the spin state |$\ket{0}$ is coherently coupled to states $\ket{+}$ and $\ket{-}$ (energetically separated by $2E$) with a Rabi frequency $\Omega$ and a detuning $\Delta$ from the average energy level of the exited states. Meanwhile, spins in states $\ket{\pm}$ feature a power-law interaction $V_{ij}=C/|i-j|^{\alpha}$. In addition to coherent drivings and interactions, each spin in states $\ket{\pm}$ decays individually to the state |$\ket{0}$ in a Markovian manner with a rate $\gamma$. The system density matrix $\rho$ then evolves according to a Lindblad master equation $\partial_{t}\rho=\mathcal{L}(\rho)$, where the Liouvillian operator acts as
@@
Mean-field analysis.—We first treat the system at the mean-field level, in which we neglect the correlation between different sites and assume that local operators $o_{i}$ have a homogeneous expectation value, i.e., $\langle o_{i}o_{j}\rangle=\langle o_{i}\rangle\langle o_{j}\rangle$ for $i\neq j$ and $\langle{o}_{i}\rangle=\langle{o}_{j}\rangle$. This treatment will generate a set of nonlinear equations [41], whose nonlinear dynamics is solely determined by the collective interaction strength $\chi=\sum_{j}V_{ij}$.
We then perform stability analysis of the equations of motion [42] and study the long-time behavior of the system. First, when states $\ket{\pm}$ are degenerate ($E=0$), the system only supports a stationary phase (SP) where observables rapidly decay to the steady-state values. Within the SP, by tuning the driving parameters $\Omega$ and $\Delta$, the system can undergo a first-order phase transition between a bistable phase and a trivial one. In fact, although $\ket{\pm}$ decay independently, the long-time behavior of the system at $E=0$ is equivalent to that of a spin-1/2 model [41], which is widely used to describe the optical bistability in Rydberg atomic ensembles [43, 44, 45, 46, 47, 48].
As the energy separation $E$ increases, the competition between states $\ket{\pm}$ comes into play, which is quantified by the order parameter $s_{z}=\langle s_{i}^{z}\rangle$ measuring their population difference. Near the regime ($\Omega\sim E$, $\Delta\sim-\chi/2$), the system supports a Hopf bifurcation at which the stability of a fixed point suddenly changes [41]. Figure 1(d) shows a typical case $E=4$, $\Delta=-7$, $\Omega=3$, $\chi=16$, $\gamma=1$ (used throughout this work), in which $s_{z}$ exhibits a persistent periodic oscillation, whose trajectory is depicted by the middle panel of Fig. 2(b). The oscillation pattern of observables is akin to the coupled pendulum model [49], with the two distinct atomic coherences acting as the harmonic oscillators. This is distinct from the limit cycle in purely dissipative cases, e.g., the Lotka–Volterra predator–prey model [50, 51, 52]. Such an understanding is confirmed by adiabatic elimination of coherence terms: in their absence, the system ceases to oscillate.
@@
We first study the Liouvillian spectrum $\{\lambda_{k}\}$ with $\mathcal{L}(\rho_{k})=\lambda_{k}\rho_{k}$ in the BTC phase. As the system size increases from 10 to 20, we identify multiple ordered branches of eigenvalues $\lambda_{k}$ shifting towards the imaginary axis, which is a universal feature for systems supporting limit cycles [57]. The inset of Fig. 3(a) depicts the low-lying spectrum for $N=20$, where we can identify a branch of eigenvalues evenly spaced along the imaginary-axis direction, as indicated by the blue dots close to the dashed grids $0,\pm\omega,\pm 2\omega,\cdots$, with $\omega$ the oscillation frequency of the limit cycle. This branch will gradually form a parabolic shape, responsible for a quantum-state diffusion process [57], and eventually collapse onto the imaginary axis in the thermodynamic limit, generating all possible oscillation patterns of the BTC. Up to $N=20$, we find that the minimum gap of this branch follows an algebraic decay $\mathrm{Re}[\lambda]\propto-N^{-\beta}$ with $\beta\approx 0.64$.
@@
In addition to the gapless branch, we find eigenvalues whose imaginary parts are incommensurate with $\omega$. The ones closest to the imaginary axis are indicated by the red diamonds in Fig. 3(a). These Liouvillian eigenstates are gapped, and determine the time scale at which a stable oscillation can be established. The gap $\approx 0.59\gamma$ drawn from an algebraic fitting shows a nice agreement with the decay rate $\approx 0.60\gamma$ toward the limit cycle.
@@
Figure 4: Scaling of the steady-state fluctuations $F^{2}$ with $N$. (a) and (b) show results for the BTC phase ($\Delta=-7$) and the SP ($\Delta=-1$), respectively. The insets show the time evolution of the fluctuations ($N=80$ for the BTC and $N=40$ for the SP), obtained from the MPS (red curves) and the CME (blue curves). The errorbars indicate the temporal fluctuations caused by finite trajectory realizations.
@@
In fact, a time crystal defined by Eq. (3) possesses not only temporal order but also long-range spatial order [67], which is indeed true for all-to-all interaction since $\mathcal{C}_{\mathrm{off}}\approx F^{2}/N=G(0)$ remains finite in the large-$N$ limit. As shown in Fig. 5(c), $\mathcal{C}_{\mathrm{off}}$ decreases monotonically with an increasing $\alpha$ and eventually vanishes when entering the SP  [^2] This trend is verified by the MPS for $N=20$ [41]. The critical property is then analyzed through finite-size scaling by assuming a universal form of $\mathcal{C}_{\mathrm{off}}=N^{\zeta/\nu}\mathcal{G}[N^{1/\nu}(\alpha-\alpha_{c})]$. We obtain a nice data collapse [see the inset of Fig. 5(c)] and extract a critical value $\alpha_{c}\approx 1.22$ together with two critical exponents $(\zeta,\nu)\approx(-0.56,1.40)$. We note that $\alpha_{c}$ is generally larger than 1 but not rigid under variation of system parameters, typical for long-range interacting systems [69].
@@
The proposed model can be potentially realized in artificial quantum systems, such as trapped ions or neutral atoms. In the former case, tunable power-law interactions with $30\leq\alpha<$ 3 can be achieved [72, 73, 74], aligning with the range considered in this work. For the latter, all-to-all interactions can be induced by cavity QED [75, 76] or Rydberg dressing in atomic ensembles [77]. Our model can also be generalized to higher dimensions, where the long-time dynamics may be accessed by improved Krylov methods [78]. In higher dimensions, dipolar ($\alpha=3$) or van der Waals ($\alpha=6$) interactions offered by Rydberg-atom arrays [79, 80, 81] may be sufficiently long-ranged to stabilize the BTC. In addition, our studies may reveal a new universality class for nonequilibrium phases [82], and open the door to exploring versatile dynamical phases, such as quasi-time crystals [83], multistability [84, 85, 86], and measurement-induced phase transition [87].
@@ Acknowledgements.
Acknowledgments.—We thank T. Pohl, F. Russo, Z.C. Zhu, X.L. Li, C. Liang, Y. Zhang, G.M. Zhang for providing helpful suggestions, and X.Y. Wen, Z. Fu, C.P. Qin for sharing computation resources. F.Y. acknowledge valuable discussions with T. Vovk, L. Sieberer, and H. Pichler. This work is supported by the Innovation Program for Quantum Science and Technology (2021ZD0302100) and by the NSFC (Grants No. 12361131576, No. 92265205). B.B. acknowledges the support from funding by the French National Research Agency (ANR) under project ANR-24-CPJ1-0150-01 and a research grant (42085) from VILLUM FONDEN. K.M. acknowledges the support from the Carlsberg Semper Ardens project QCooL. Work in Innsbruck is supported by the ERC Starting grant QARA (Grant No. 101041435). Note added: During preparation of this manuscript, we became aware of a related work on three-level dissipative time crystals induced by a dipolar interaction in 2D [88].